OpenStudy (garypowell):

f(x) = 4x + 7, g(x) = 3x2 Find (f + g)(x). (1 points)

OpenStudy (cashout_yio):

42x^3 + 42x^2

OpenStudy (kittiwitti1):

No direct answering ! ^_^ (f+g)(x) is also written as f(x) + g(x).

OpenStudy (3mar):

you mean\[f(x)=4x+7\] \[g(x)=3x^2\]

OpenStudy (retireed):

It is similar to this .... f(x) = 3 g(x) = 10 f(x) + g(x) = 3 +10 (f+g)x = 13 Apply this logic to your problem

OpenStudy (kittiwitti1):

This gets you:\[f(x)+g(x)=(4x+7)+(3x^{2})\]

OpenStudy (kittiwitti1):

@retirEEd has the right idea.

OpenStudy (3mar):

(f+g)(x)=f(x)+g(x)

OpenStudy (3mar):

This leads to \[(f+g)(x)=4x+7+3x^2=3x^2+4x+7\]